ABC corp acquired a van on January 1, 2016 for $330,000. The company used the sum or years digits method assuming a 4-year useful life and a $10,000 residual value. During 2018 the company decides to change to straight-line depreciation and determines that the useful life of the truck is 6 years (no change to residual value).

a) what depreciation expense should ABC corp recognize in 2018?

b)how will these changes affect net income in 2018 assuming ABC corp pays a 20% tax rate?
Answer #1
Sum of years digits Method
Life in years                         4.00
Cost of Asset            330,000.00
Salavge Value              10,000.00
Depreciable Value =              320,000.00
sum of the years' digits depreciation calculation is = n(n+1)/2 = 4(4+1)/2 = 4*5/2 = 20/2 = 10
Year Depreciation WN
                                                 2,016.00            128,000.00 320,000/10 * 4
                                                 2,017.00              96,000.00 320,000/10 * 3
           224,000.00
Balance at beginning of 2018 = 330,000 - 224,000            106,000.00
Remaining life = 6years -2 Years                         4.00
Salavge Value              10,000.00
Depreciation expense in 2018 = (106,000 - 10,000)/4              24,000.00
If used sum of years method then dep would be 320,000/10*2              64,000.00
Difference in depreciation = 64000 - 24000              40,000.00
Difference in depreciation net of tax = 40000*.80              32,000.00
Net income would increase by 32000 because of using SLM method